# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v1

@@ III.2 Mixed Strategy
Now we recall the well-known result Schuster et al. (1981); Selten (1980) that ESS of Definition 1c or 2c can never be mixed. There may, however, be mixed 2ESS—the ESS of Definition 3c. Since 2ESS must be Nash equilibrium which must be a fixed point of the replicator equation, let us assume that it correspond to a fixed point $(\hat{x},\hat{y})$. A glance at replicator equation (21) reveals that with $\alpha_{1}\equiv\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}-\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$,  $\beta_{1}\equiv\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}-\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$,  $\alpha_{2}\equiv\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}-\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$ and $\beta_{2}\equiv\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}-\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$, the internal fixed point must be $(\hat{x},\hat{y})=\left({\alpha_{2}}/({\alpha_{2}+\beta_{2}}),{\alpha_{1}}/({\alpha_{1}+\beta_{1}})\right)$. The eigenvalues of Jacobian in linearization about this fixed points can be calculated to be ±$\pm\sqrt{({\alpha_{1}\beta_{1}\alpha_{2}\beta_{2}})/[{(\alpha_{1}+\beta_{1})(\alpha_{2}+\beta_{2})}]}$. Evidently, the fixed point can either be a center or a saddle. We know that (linear) center found using the linear stability analysis may not actually be a robust nonlinear center. However, it is known that the system is Hamiltonian Hofbauer and Sigmund (1998) in the interior of the phase space, and hence any center in this case is, in fact, nonlinear center. Thus, ESS is not strongly stable.
Our final conclusion is that Theorem 1 is incorrect if the ESS is pure and is that of Definition 3b but is correct if the pure ESS is that of Definition 1b or 2b; the theorem is incorrect even if the ESS is mixed and is that of Definition 3b. It is now natural to ask if introducing some intraspecific interactions render Theorem 1—with replicator equation modified to represent the resultant asymmetric game—true for 2ESS, i.e., the ESS of Definition 3b.
IV Effect of intraspecific interaction: Asymmetric games
@@
Due to the presence of intraspecific interaction, we must introduce two additional square payoff matrices—denoted as $\sf{A}$ and $\sf{D}$, respectively of sizes $m$ and $n$—accounting for the payoff gained by players of population X and players of population Y, respectively. The new payoff components are just the ones that one uses very commonly in symmetric games Smith (1982). Furthermore, rather than working with Definitions 1b, 2b, and 3b, it turns out to be more convenient to work equivalently with Definitions 1c, 2c, and 3c, which however now must be modified to include the intraspecific interactions. Consequently, we write the extensions—Definitions 1d, 2d, and 3d—of Definitions 1c, 2c, and 3c, respectively, as follows:
@@
  1. Definition 1d: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}\text{and}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (23)
@@
    Taylor Taylor (1979), however, proposes Taylor (1979) that for $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ to be ESS the total payoff of the resident against the mutant must be higher than that of mutant’s. Mathematically, we may write,
@@
  2. Definition 2d: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
@@
  3. Definition 3d: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\text{either}~{}~{}\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}\text{or}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (25)
    ESS under this definition is also called 2ESS as before.
@@
  1. Theorem ${\bm{1}^{\prime}}$: A 2ESS is a locally asymptotically stable rest point of the following replicator equations,
@@ IV.1 Pure ESS
Now, just like we have argued in the last section rather elaborately using the concept of strong stability and illustrative counterexample that Theorem 1 is correct only when the 2ESS in its statement is replaced by ESS of Definitions 1c or 2c, here we want to argue that Theorem 1 is also not valid for pure ESS unless 2ESS in its statement is replaced by the ESS of Definition 1d or Definition 2d. We achieve this goal in two ways. First, we demonstrate that up to linear order, an ESS that satisfies either Definition 1d or Definition 2d is linearly stable. In contrast, an ESS that satisfies Definition 3d but not 1d or 2d is not linearly stable. Next, through the recognition of a Lyapunov function, we show that if either of Definitions 1d or 2d is satisfied, the corresponding fixed point is locally asymptotically stable.
@@
To build on this, we extend our approach beyond the restriction of two strategies per subpopulation and address the issue in its full generality. We remind ourselves that these definitions and the theorem are valid for any finite number of strategies, i.e., they are valid for arbitrary $m$ and $n$ values. In fact, even though our focus here is on asymmetric games, it formally extends our conclusion on bimatrix games for arbitrary $m$ and $n$ values—all one has to do is set ${\sf A}={\sf D}={\sf 0}$ in what follows.
@@
Let $a_{ij}$’s, $b_{ij}$’s, $c_{ij}$’s and $d_{ij}$’s denote the elements of the payoff matrices ${\sf A}$, ${\sf B}$, ${\sf C}$ and ${\sf D}$, respectively. Without any loss of generality, consider the state under investigation be $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ which is a fixed point of the set of equations given by Eqs. (26a) and (26b). On doing linear stability analysis about it yields a diagonal Jacobian matrix, and hence the diagonal elements can be easily read as the eigenvalues—say, $\lambda_{1},\lambda_{2},\cdots,\lambda_{m-1},\lambda^{\prime}_{1},\lambda^{\prime}_{2},\cdots,\lambda^{\prime}_{n-1}$—which respectively are $a_{1m}+b_{1n}-a_{mm}-b_{mn},a_{2m}+b_{2n}-a_{mm}-b_{mn},\cdots,a_{m-1,m}+b_{m-1,n}-a_{mm}-b_{mn},c_{1m}+d_{1n}-c_{nm}-d_{nn},c_{2m}+d_{2n}-c_{nm}-d_{nn},\cdots,c_{n-1,m}+d_{n-1,n}-c_{nm}-d_{nn}$. Here, comma has been put between subscripts wherever it is needed for clarity.
@@
Now for finding ESS, we write the arbitrary vector in the neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ as $({\mathbf{x}},{\mathbf{y}})=((\epsilon_{x_{1}},\cdots,\epsilon_{x_{m-1}},1-\epsilon_{x_{1}}-\cdots-\epsilon_{x_{m-1}}),(\epsilon_{y_{1}},\cdots,\epsilon_{y_{n-1}},1-\epsilon_{y_{1}}-\cdots-\epsilon_{y_{n-1}}))$, where the $\epsilon$’s are infinitesimal. A glance at Definition 1d directs us to calculate following terms, which we do up to first order in $\epsilon$’s:
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{m1}-a_{mm})\epsilon_{x_{1}}+(a_{m2}-a_{mm})\epsilon_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\epsilon_{x_{m-1}},$ (27a)
$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\epsilon_{y_{1}}+(b_{m2}-b_{mn})\epsilon_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\epsilon_{y_{n-1}},$ (27b)
$\mathbf{x}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{1m}-a_{mm})\epsilon_{x_{1}}+(a_{2m}-a_{mm})\epsilon_{x_{2}}+\cdots+(a_{m-1,m}-a_{mm})\epsilon_{x_{m-1}}$ (27c)
$~{}~{}~{}~{}~{}+(a_{m1}-a_{mm})\epsilon_{x_{1}}+(a_{m2}-a_{mm})\epsilon_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\epsilon_{x_{m-1}},$
$\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\epsilon_{y_{1}}+(b_{m2}-b_{mn})\epsilon_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\epsilon_{y_{n-1}}$ (27d)
$~{}~{}~{}~{}~{}+(b_{1n}-b_{mn})\epsilon_{x_{1}}+(b_{2n}-b_{mn})\epsilon_{x_{2}}+\cdots+(b_{m-1,n}-b_{mn})\epsilon_{x_{m-1}}.$
@@
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=-\epsilon_{x_{1}}\lambda_{1}-\epsilon_{x_{2}}\lambda_{2}-\cdots-\epsilon_{x_{m-1}}\lambda_{m-1}.$ (28)

### v2

@@ III.3 Mixed ESS
Now we recall the well-known result Schuster et al. (1981); Selten (1980) that ESS of Definitions 1c or 2c can never be mixed. There may, however, be mixed 2ESS—the ESS of Definition 3c (as shown below) under simultaneous deviation perspective. Since 2ESS must be NE which must be a fixed point of the replicator equation, let us assume that it corresponds to a fixed point $(\hat{x},\hat{y})$. A glance at replicator equation (19) reveals that with $\alpha_{1}\equiv\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}-\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$,  $\beta_{1}\equiv\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}-\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$,  $\alpha_{2}\equiv\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}-\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$ and $\beta_{2}\equiv\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}-\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$, the internal fixed point must be $(\hat{x},\hat{y})=\left({\alpha_{2}}/({\alpha_{2}+\beta_{2}}),{\alpha_{1}}/({\alpha_{1}+\beta_{1}})\right)$. The eigenvalues of Jacobian in linearization about this fixed points can be calculated to be ±$\pm\sqrt{({\alpha_{1}\beta_{1}\alpha_{2}\beta_{2}})/[{(\alpha_{1}+\beta_{1})(\alpha_{2}+\beta_{2})}]}$. Evidently, the fixed point can either be a center or a saddle. We know that (linear) center found using the linear stability analysis may not actually be a robust nonlinear center. However, it is known that the system is Hamiltonian Hofbauer and Sigmund (1998) in the interior of the phase space, and hence any center in this case is, in fact, nonlinear center. Thus, ESS of any definition is not strongly stable.
In conclusion, we find that Theorem 1 is invalid if the 2ESS is mixed. The theorem, however, can be saved by either adopting face-value perspective which allows for no mixed ESS or, even more interestingly, using another version of the replicator equation, viz., adjusted replicator equation Mukhopadhyay and Chakraborty (2021); Smith (1982); Hofbauer and Sigmund (1998):
$\dot{x}_{i}=x_{i}\frac{({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf B}\mathbf{y}}{\mathbf{x}\cdot{\sf B}\mathbf{y}},$ (29a)
$\dot{y}_{i}=y_{i}\frac{({\sf C}\mathbf{x})_{i}-\mathbf{y}\cdot{\sf C}\mathbf{x}}{\mathbf{y}\cdot{\sf C}\mathbf{x}}.$ (29b)
Under its dynamics, the mixed state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ can be asymptotically stable (see Fig. 1c).
Example 3 (see Fig. 1c): Mixed 2ESS under simultaneous deviation perspective is center-type fixed point in $2\times 2$ bimatrix games.
To simplify the calculation while keeping generality, we use two key facts: first, one can always modify the payoff matrix by adding a matrix with identical rows and by scaling the payoff matrices, which preserves both the dynamics and the ESS. Specifically, ${\sf B}$ and $c{\sf B}+{\sf\Omega}$ are equivalent, where $c$ is a positive scalar and ${\sf\Omega}_{ij}=\omega_{j}$. Second, using this transformation, one can show that any $2\times 2$ bimatrix game can be reduced to either a rescaled zero-sum game or a rescaled partnership game Hofbauer and Sigmund (1998). Therefore, we can work with the following payoff matrix:
${\sf B}=\begin{pmatrix}\beta_{1}&0\\ 0&\alpha_{1}\end{pmatrix}~{}\text{and}~{}{\sf C}=\begin{pmatrix}\beta_{2}&0\\ 0&\alpha_{2}\end{pmatrix},$ (30)
where for rescaled partnership game, $\alpha_{1}=c\alpha_{2}$ and $\beta_{1}=c\beta_{2}$, on the other hand for rescaled zero-sum game $\alpha_{1}=-c\alpha_{2}$ and $\beta_{1}=-c\beta_{2}$ Hofbauer and Sigmund (1998). Now, the internal fixed point—if it exist in the simplotope—is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}\alpha_{2}/(\alpha_{2}+\beta_{2}),{\beta_{2}}/({\alpha_{2}+\beta_{2}})\end{pmatrix},\begin{pmatrix}{\alpha_{1}}/({\alpha_{1}+\beta_{1}}),{\beta_{1}}/({\alpha_{1}+\beta_{1}})\end{pmatrix}\right)$ and the neighbourhood vector is $(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}\alpha_{2}/(\alpha_{2}+\beta_{2})+\delta_{x},{\beta_{2}}/({\alpha_{2}+\beta_{2}})-\delta_{x}\end{pmatrix},\begin{pmatrix}{\alpha_{1}}/({\alpha_{1}+\beta_{1}})+\delta_{y},{\beta_{1}}/({\alpha_{1}+\beta_{1}})-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small which can have positive or negative sign but neither of them are allowed to be zero as per simultaneous deviation perspective. We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{y}=-\delta_{x}\delta_{y}(\alpha_{1}+\beta_{1})$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=-\delta_{x}\delta_{y}(\alpha_{2}+\beta_{2})$. We can see that the fixed point cannot be an ESS under or-definition if $\alpha_{1}+\beta_{1}$ and $\alpha_{2}+\beta_{2}$ have the same sign, which is the case of rescaled partnership game. The fixed point is always an ESS under or-definition if the they have different sign, this is the case of rescaled zero-sum game. Next we note that the eigenvalues (of Jacobian during linearization about the fixed point) simplify to ±$\pm\sqrt{c}({\alpha_{2}\beta_{2}/(\alpha_{2}+\beta_{2})})$ for rescaled partnership game and ±$\pm i\sqrt{c}({\alpha_{2}\beta_{2}/(\alpha_{2}+\beta_{2})})$ for rescaled zero-sum game, which verifies the claim of ESS being a center.
Therefore, we conclude that mixed ESS under or-definition is possible, however it does not tell anything about asymptotic stability—an ESS can at best be Lyapunov stable as a center. This also tells us that unlike the case of strict NE, and-, sum- and or-definitions are not equivalent when it comes to mixed NE; in fact there is no scope of mixed ESS under and- and sum-definitions. Finally, of course, as per the face-value perspective the interior fixed point is not an ESS, since $\delta_{x}=0$ makes $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=0$.
IV Asymmetric games
It is now natural to ask if introducing some intra-specific interactions changes this situation, i.e., renders Theorem 1—with replicator equation modified to represent the resultant asymmetric game—true for a mixed ESS and makes it possible for a mixed NE to be an ESS under and- and sum-definitions (modified appropriately for asymmetric games). In the present section, we will discuss how adding intra-specific interactions addresses these questions and makes the concept of 2ESS more useful in the context of asymmetric games (see also Appendix A).
@@ IV Asymmetric games
Due to the presence of intra-specific interaction, we must introduce two additional square payoff matrices—denoted as $\sf{A}$ and $\sf{D}$, respectively of sizes $m$ and $n$—accounting for the payoff gained by players of population X and players of population Y, respectively. The new payoff components are just the ones that one uses very commonly in symmetric games Smith (1982). Furthermore, rather than working with Definitions 1b, 2b, and 3b, it turns out to be more convenient to work equivalently with Definitions 1c, 2c, and 3c, which however now must be modified to include the intra-specific interactions. Consequently, we write the extensions—Definitions 1d, 2d, and 3d—of Definitions 1c, 2c, and 3c, respectively, as follows:
@@
  1. Definition 1d (and-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}\text{\emph{and}}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (31)
@@
    Taylor Taylor (1979), however, proposes that for $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ to be ESS the total payoff of the resident against the mutant must be higher than that of mutant’s. Mathematically, we may write,
@@
  2. Definition 2d (sum-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
@@
  3. Definition 3d (or-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\text{either}~{}~{}\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}\text{\emph{or}}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (33)
    ESS under this definition is also called 2ESS as before. When both inter- and intra-specific interactions are simultaneously present, mixed 2ESS may appear even without the imposing of additional simultaneous deviation perspective, which thus is no longer adopted henceforth.
@@
  1. Theorem ${\bm{1}^{\prime}}$: An ESS is a locally asymptotically stable rest point of the following replicator equations,
@@
We now turn to discussing pure and mixed ESS’s separately in the context of asymmetric games.
@@ IV.1 Pure ESS
In case of a strict NE being ESS, all the eigenvalues of the Jacobian about the corresponding fixed point of the replicator dynamics are negative making the fixed point locally asymptotically stable under all the three definitions (Definitions 1d, 2d, and 3d). To demonstrate this, we go beyond the restriction of two strategies per subpopulation and address the matter in its full generality. We remind ourselves that these definitions are valid for any finite number of strategies, i.e., for arbitrary $m$ and $n$ values. In fact—as mentioned in Section III.2—even though our focus here is on asymmetric games, it formally extends our conclusion on bimatrix games for arbitrary $m$ and $n$ values—all one has to do is set ${\sf A}={\sf D}={\sf 0}$ in what follows.
IV.1.1 Strict NE as ESS
@@
Let $a_{ij}$’s, $b_{ij}$’s, $c_{ij}$’s and $d_{ij}$’s denote the elements of the payoff matrices ${\sf A}$, ${\sf B}$, ${\sf C}$ and ${\sf D}$, respectively. Without any loss of generality, consider the state under investigation be $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ which is a fixed point of the set of equations given by Eqs. (34a) and (34b). On doing linear stability analysis about it yields a diagonal Jacobian matrix, and hence the diagonal elements can be easily read as the eigenvalues—say, $\lambda_{1},\lambda_{2},\cdots,\lambda_{m-1},\lambda^{\prime}_{1},\lambda^{\prime}_{2},\cdots,\lambda^{\prime}_{n-1}$—which respectively are $a_{1m}+b_{1n}-a_{mm}-b_{mn},a_{2m}+b_{2n}-a_{mm}-b_{mn},\cdots,a_{m-1,m}+b_{m-1,n}-a_{mm}-b_{mn},c_{1m}+d_{1n}-c_{nm}-d_{nn},c_{2m}+d_{2n}-c_{nm}-d_{nn},\cdots,c_{n-1,m}+d_{n-1,n}-c_{nm}-d_{nn}$. Here, comma has been put between subscripts wherever it is needed for clarity. It is clear that all the eigenvalues must be negative if the fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ is a strict NE; thus, linear stability is automatically implied in such cases.
@@
Now we would like to figure out whether the strict NE is ESS. For this purpose, we write the arbitrary vector in the neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ as $({\mathbf{x}},{\mathbf{y}})=((\delta_{x_{1}},\cdots,\delta_{x_{m-1}},1-\delta_{x_{1}}-\cdots-\delta_{x_{m-1}}),(\delta_{y_{1}},\cdots,\delta_{y_{n-1}},1-\delta_{y_{1}}-\cdots-\delta_{y_{n-1}}))$, where the $\epsilon$’s are infinitesimal and non-negative. A glance at Definition 1d directs us to calculate following terms, which we do up to first order in $\delta$’s:
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{m1}-a_{mm})\delta_{x_{1}}+(a_{m2}-a_{mm})\delta_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\delta_{x_{m-1}},$ (35a)
$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\delta_{y_{1}}+(b_{m2}-b_{mn})\delta_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\delta_{y_{n-1}},$ (35b)
$\mathbf{x}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{1m}-a_{mm})\delta_{x_{1}}+(a_{2m}-a_{mm})\delta_{x_{2}}+\cdots+(a_{m-1,m}-a_{mm})\delta_{x_{m-1}}$ (35c)
$~{}~{}~{}~{}~{}+(a_{m1}-a_{mm})\delta_{x_{1}}+(a_{m2}-a_{mm})\delta_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\delta_{x_{m-1}},$
$\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\delta_{y_{1}}+(b_{m2}-b_{mn})\delta_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\delta_{y_{n-1}}$ (35d)
$~{}~{}~{}~{}~{}+(b_{1n}-b_{mn})\delta_{x_{1}}+(b_{2n}-b_{mn})\delta_{x_{2}}+\cdots+(b_{m-1,n}-b_{mn})\delta_{x_{m-1}}.$
@@
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=-\delta_{x_{1}}\lambda_{1}-\delta_{x_{2}}\lambda_{2}-\cdots-\delta_{x_{m-1}}\lambda_{m-1}.$ (36)